# Nonlinear Fore(Back)casting and Innovation Filtering for Causal-Noncausal VAR Models

### v3

@@ 5.3 Confidence Interval of the Prediction Interval
The analysis of a confidence set of the prediction set is related to research on a confidence set of the identified set in models under partial identification. The partial identification literature considers a parametric model with a parameter $\gamma$, say, that is partly identifiable  [^11] See Imbens, Manski (2004) for confidence intervals of identified intervals in the framework of partial identification and confidence intervals that asymptotically cover the true interval with a probability larger or equal to $1-\alpha_{2}$.. The objective is to determine either the confidence set under the classical approach, or the credible set under the Bayesian approach. In our framework, we have two types of ”parameters”, $P$ and $Y_{1,T+1}$, say. The second one is not identifiable, although its conditional distribution is estimated, and plays the role of a conditional prior.
@@ 6 Illustration
We illustrate the nonlinear forecasts from the mixed VAR(1) model. The first part of this Section presents a simulation study that examines the out of sample forecasts from the model. Next, the semi-parametric GCov estimation, nonlinear innovations and CBS are illustrated in applications to the joint analysis of a bivariate series of Bitcoin/USD and Ethereum/USD exchange rates.
@@ 6.1 Simulation Study
6.1.1 The Artificial Data Set
We consider a simulated bivariate mixed VAR(1) process with the following matrix $\Phi$ of autoregressive coefficients and matrix A:
$\Phi=\left(\begin{array}[]{cc}0.7&-1.3\\ 0&2\end{array}\right),\;\;A=\left(\begin{array}[]{cc}1&-1\\ 0&1\end{array}\right).$
where matrix $\Phi$ has eigenvalues 0.7 and 2, located inside and outside the unit circle, respectively  [^12] The finite sample properties of the GCov estimator are illustrated in Gourieroux, Jasiak (2017).. The errors follow a bivariate noise with independent components both t-student distributed with $\nu$ =4 degrees of freedom, mean zero and variance equal to $\nu/(\nu-2)=2$.
The simulated paths of the series of length 600 are displayed in Figure 1. The solid (black) line represents process $Y_{1t}$ and the dashed (red) line represents process $Y_{2t}$.
@@ 6.1.1 The Artificial Data Set
The sequence of spikes in the noncausal component $Y_{2t}=Z_{2t}$ impacts the component $Y_{1t}$ through the recursive form of matrix $\Phi$.
6.1.2 Estimated Predictive Density, Point and Interval Forecasts
Let us now consider forecasts based on the estimated model parameters. The Generalized Covariance (GCov) estimate of matrix $\Phi$ is obtained by minimizing the portmanteau statistic computed from the auto- and cross-correlations up to and including lag $H=10$ of the errors $\varepsilon_{t}=Y_{t}-\Phi_{1}Y_{t-1}$ and their squared values [see Gourieroux, Jasiak (2022)]. The following estimated autoregressive matrix is obtained:
$\hat{\Phi}=\left(\begin{array}[]{cc}0.724&-1.452\\ -0.030&1.993\end{array}\right),$
with eigenvalues $\hat{\lambda}_{1}=0.690$, $\hat{\lambda}_{2}=2.027$ close to the true values $\lambda_{1}=J_{1}=0.7$ and $\lambda_{2}=J_{2}=2.0$. The standard errors of $\hat{\Phi}$ obtained by bootstrap are 0.023, 0.308, for the elements of the first row, and 0.009, 0.120 for the elements of the second row.
@@ 6.1.2 Estimated Predictive Density, Point and Interval Forecasts
After estimating $\Phi$, the GCov estimated errors $\hat{\varepsilon}_{t}=Y_{t}-\hat{\Phi}Y_{t-1}$ are computed and used for forecasting. Matrices $A$ and $A^{-1}$ are identified from the Jordan representation of matrix $\Phi$, up to scale factors. The estimated matrix $\hat{A}^{-1}$ computed from the normalized Jordan decomposition of $\hat{\Phi}$ is:
@@
$\hat{A}^{-1}=\left(\begin{array}[]{cc}0.022&0.025\\ -0.022&0.974\end{array}\right)$. It corresponds to matrix $A^{-1}=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)$ up to scale factors of about 0.02 and 0.97 for each column.

### v4

@@ 4.3 Confidence Set of the Prediction Interval
The analysis of confidence set for the prediction interval is related to set identification, where a confidence set for the identified set is determined for models under partial identification [Beresteanu et al., (2017)]. The partial identification literature considers a parametric model with a partly identifiable parameter [^7] See Imbens and Manski (2004) for confidence intervals of identified intervals in the framework of partial identification and confidence intervals that asymptotically cover the true interval with a probability larger or equal to $1-\alpha_{2}$. to determine either the confidence set under the classical approach, or the credible set under the Bayesian approach. In our framework, we have two types of ”parameters”: $P$ and $Y_{1,T+1}$. The second one is not identifiable, although its conditional distribution is estimated, and plays the role of a conditional prior.
5 Nonlinear Causal Innovations
The inference on the traditional causal VAR model commonly includes the IRF analysis, which is an important tool used by economists and financial policy makers for the so-called causal analysis. This terminology differs from the ”causal-noncausal” terminology introduced by Rosenblatt for time series. More precisely, the ”causal analysis” involves a) making inference on the future given the past, i.e. forecasting, and also b) the counterfactual analysis of the effects of current transitory shocks on the future. The standard linear causal IRF analysis cannot be applied to the mixed VAR(p) processes, because neither errors $\varepsilon_{t}$ in model (2.1), nor the state-specific linear errors $\eta_{t}$ in (2.5) are causal and independent of the lagged values of $Y_{t}$. Hence, it is difficult to interpret the shock on $\varepsilon_{t}$ that implicitly changes a past that has already been realized. This section introduces an alternative concept of innovation that eliminates such a difficulty in nonlinear IRF analysis. This new innovation accounts for nonlinear dynamics of $(Y_{t})$ with bubbles and spikes. In addition, it satisfies both the serial and cross-sectional independence conditions [Gourieroux and Jasiak (2005), Gourieroux, Monfort and Renne (2017) on structural SVAR models in Macroeconomics]. We discuss below the filtering and identification of the nonlinear causal innovations in mixed VAR models.
5.1 Definition of Nonlinear Causal Innovations
Definition 1: Let us consider a Markov process of order $p$ with a positive continuous transition density. A nonlinear causal innovation of process $(Y_{t})$ is a process $(v_{t})$ of dimension $m$ such that: i) the vectors $v_{t}$ are serially i.i.d.; ii) the strictly stationary process $(Y_{t})$ can be written in a nonlinear causal autoregressive form:
$Y_{t}=a(\underline{Y_{t-1}},v_{t}),\;\forall t,$ (5.1)
with $\underline{Y_{t-1}}=(Y_{t-1},...,Y_{t-p})$; iii) the future values $\bar{v}_{t}$ of the innovation process are independent of the lagged values $\underline{Y_{t-1}}$ of the observed process $(Y_{t})$.
It is easy to see that conditions i) and ii) are equivalent to the Markov of order $p$ property of $(Y_{t})$ with a continuous distribution, and can be applied in particular to the mixed model (2.1) by Corollary 1. Such a nonlinear causal autoregressive form always exists [see Rosenblatt (1952) and the discussions in Sections 5.2, 5.3]. It is not unique in a nonparametric framework since $v_{t}$ can always be replaced by $w_{t}=d(v_{t})$, where $d$ is a diffeomorphism, and $a$ is replaced by $a\circ d^{-1}$. If function $a$ is invertible with respect to $v$, then $v_{t}$ is a nonlinear function of $Y_{t}$ given its past and condition iii) is satisfied. Under conditions i) and ii) the autoregressive equation (5.1) can be applied recursively.
The sequence of nonlinear causal innovations provides the basis of nonlinear impulse response functions [Gourieroux and Jasiak (2005), Gonzalves et al. (2021)]. Let us consider a transitory shock $\delta$ at date $T$ on $v_{T}$, and then apply recursively the autoregressive equation (5.1) to get the IRF, defined as shocked future values $Y_{T}^{\delta}=a(Y_{T-1},Y_{T-2},\cdots,v_{T}+\delta)$, $Y_{T+1}^{\delta}=a(Y_{T}^{\delta},Y_{T-1},...,v_{T+1})$, and so on. The conditions i) and iii) in Definition 1 are crucial for the interpretation of these shocks in terms of ”causal analysis”. Condition iii) means that $v_{T}$ can be shocked at any time T without an effect on the realized past $\underline{Y}_{T-1}$. Condition i) means that this shock has no effect on the future values $\bar{v}_{T+1}=(v_{T+1},v_{T+2},...)$.
The nonlinear IRF defined above corresponds to a multivariate shock $\delta$. The literature on ”causal analysis” is also interested in univariate shocks to specific variables. It is possible to define shocks to a specific component $v_{1,t}$ of $v_{t}$, if additionally the component $v_{1,t}$ and ($v_{2,t},...,v_{n,t}$) are cross-sectionally independent.
5.2 Identification of Nonlinear Autoregression and Innovation
In the linear causal VAR model, there are identification problems concerning the parameters of the model and the innovation-based IRFs, because the innovations to be shocked are defined up to an orthogonal linear transform. Both these identification problems are solved by identifying the mixing matrix $D$ in the $\varepsilon_{t}=Dv_{t}$ error representation, using the Independent Component Analysis (ICA) method for example [see Gourieroux, Monfort and Renne (2017)].
The nonlinear autoregression (5.1) depends on function $a$ and the distribution of the causal nonlinear innovation $v_{t}$. In the mixed VAR(p) model, both depend on $\Phi_{1},\Phi_{2},...$ and $g$, without being one-to-one functions of these parameters, because the nonlinear function $a$ depends on more arguments than the density $g$. This new nonparametric identification issue concerning the causal innovation can be examined in a functional framework by using the properties of harmonic functions. In particular, we get the following result:
Proposition 2 [Gourieroux and Lee (2025)]: In the nonlinear causal autoregressive representation of a mixed VAR(1) process, the dimension of under-identification in the functional space of nonlinear autoregressive models is finite and equal to $2m$.
This result reveals the challenges and complexity of the identification of the nonlinear causal autoregression. The above under-identification is very large and exceeds the under-identification of linear causal VAR models. As pointed out in Section 5.1, it is possible to restrict the analysis to nonlinear autoregressions with Gaussian innovations, such that the components $v_{1,t},...,v_{n,t}$ are independent N(0,1). However, this additional restriction is insufficient to solve the identification issue. Indeed, there exist nonlinear transformations of $v_{t}$, with respect to which the multivariate N(0, Id) distribution remains invariant [see Gourieroux and Lee (2025)], corresponding to ”local matrix rotations”.
Corollary 3: Functional multiplicity There exists a functional multiplicity of innovation processes ($v_{t}$), which are cross-sectionally independent standard normal variables.
It follows that the functional identification of the nonlinear autoregressive representation of a causal nonlinear innovation and nonlinear IRFs is a complicated issue in nonlinear autoregressive models, and in particular in the mixed VAR models. It resembles the identification issue of causal VAR and Structural (SVAR) models, which has remained unsolved for a long time. Until a better solution becomes available, we can use one of the two following methods:
i) We can introduce identifying restrictions on function $a$ through parametric assumptions on the joint distribution of errors $\varepsilon_{t}$. For example, one could assume that this distribution is a non-Gaussian elliptical or t-Student. However, such a partial identifying assumption is difficult to justify due to the lack of a structural interpretation of errors ($\varepsilon_{t}$).
ii) A recursive structure of the model can be assumed, by analogy to the recursive structure of errors in causal VAR models that has been traditionally used for (S)VAR identification in the early macroeconomic literature [see Sims (1980) based on Cholesky decomposition].
5.3 Shock Ordering
In practice one may prefer to adopt method ii). Note that since the observed process ($Y_{t}$) and the state process ($Z_{t}$) satisfy a one-to-one relationship, the state variables also satisfy a nonlinear autoregressive scheme:
$Z_{t}=\tilde{a}(\underline{Z}_{t-1},\tilde{v}_{t}),$ (5.2)
where $\tilde{v}_{t}$ can be serially conditionally i.i.d. (and possibly restricted to be standard Gaussian). A recursive structure can be imposed on model (5.2) with the noncausal state variable to be shocked first under this approach (also referred to as ”shock ordering”). This choice can be further motivated by the fact that in applied research, the noncausal order is often equal to 1, i.e. $n_{2}=1$ [see e.g. Hecq, Lieb and Telg (2016), Gourieroux and Jasiak (2017) and Section 6.3]. This implies that there is a common noncausal component generating the bubbles and local trends in $(Y_{t})$. For example, in macroeconomic models, that noncausal component can be related to speculative bubbles in oil prices impacting jointly the price index, GDP and other macroeconomic variables. In financial applications, the single factor capturing the nonlinear dynamics can be interpreted as a systemic component, or a common bubble [Gourieroux and Zakoian (2017), Cubadda et al. (2023), Hall and Jasiak (2024)]. For a bivariate VAR(1) with $n_{2}=1$, we get an identifiable recursive form of the nonlinear autoregressive model (5.2):
$Z_{2,t}$ $=$ $G_{2}(Z_{2,t-1};v_{2,t}),$ (5.3)
$Z_{1,t}$ $=$ $G_{1|2}(Z_{2,t},Z_{1,t-1};v_{1,t}),$ (5.4)
where function $G_{2}$ in $Z_{2,t}=G_{2}(Z_{2,t-1};v_{2,t})$ is the inverse of the function defining the nonlinear Gaussian causal innovation of $Z_{2}$, which is uniquely defined as:
$v_{2,t}(Z)=\Phi^{-1}[F_{2}(Z_{2,t}|\underline{Y}_{t-1})]=\Phi^{-1}[F_{2}(Z_{2,t}|Z_{t-1})],$ (5.5)
where $F_{2}$ is the conditional cumulative distribution function (c.d.f) of $Z_{2,t}$ given $Z_{t-1}$ and $\Phi$ is the c.d.f. of the standard Gaussian distribution. From equation (a.5) given in Appendix A.1, it follows that the conditional density of $Z_{2,t}$ given $Z_{t-1}$ has a closed form given by:
where $g_{\eta_{2}}$ is the marginal density of $\eta_{2}$. The c.d.f. $F_{2}$ is the integral of $l(z_{2,t}|z_{t-1})$ over the set of admissible values of $z_{2,t}$. By the Markov property of $(Z_{2,t})$ in Corollary 4 below, $Z_{1,t-1}$ does not appear in equation (5.6) and we have $v_{2,t}=\Phi^{-1}[F_{2}(Z_{2,t}|Z_{2,t-1})]$.
Next, we append $v_{2,t}$ by the (Gaussian) ($v_{1,t}$) associated with the causal state variable(s) in this recursive model. In our illustration, we consider the case $n_{1}=1$ and define:
$v_{1,t}(Z)=\Phi^{-1}[F_{1|2}(Z_{1,t}|Z_{2,t},Z_{t-1})],$ (5.7)
where $F_{1|2}$ is the conditional cumulative distribution function of $Z_{1,t}$ given $Z_{2,t},Z_{t-1}$  [^8] The system of equations (5.6)-(5.8) is the inverse of system (5.5)-(5.7) and it can be used for simulating the path of $Z_{t}$ from independent standard Gaussian drawing of $v_{1,t},v_{2,t}$.. From equation (a.5) given in Appendix A.1, we get the closed-form expression of the conditional density:
When the noncausal state variable $Z_{2}$ is the first one to be shocked, it is easy to check that $v_{1,t}(Z)$ differs in general from the causal innovation $\eta_{1,t}$ associated with the latent causal component $Z_{1,t}$. By applying the shock ordering, we are implementing the autoregressive model (5.2) in a recursive form (5.3)-(5.4). Moreover, we deduce from (5.5):
Corollary 4: The noncausal state variable $(Z_{2,t})$ is a Markov process of order 1 with respect to the filtration associated with $(Y_{t})$, [and also to the filtrations associated with $(Z_{t})=(Z_{1,t},Z_{2,t})$, and with $(Z_{2,t})]$.
By construction, the structural noncausal shock $v_{2,t}(Z)$ is defined in a unique way (since $n_{2}=1$) and is a Gaussian white noise [^9] See Gourieroux and Jasiak (2005) for the uniqueness of a nonlinear Gaussian innovation in univariate models.. It is also independent of the multivariate shock $v_{1,t}(Z)$. Then, we can trace out nonlinear IRFs based on the effect of a change $\delta_{2}$ in $v_{2,t}$, with $v_{1,t}$ held constant. It is called henceforth the Common Bubble Shock (CBS) [See Online Appendix C for quantile estimation of $F_{2}(Z_{2,t}|Z_{2,t-1})$]. In the general case $n_{2}=1,n_{1}\geq 1$, the associated multivariate impulse responses of $Z_{2,t}$ are identifiable, i.e. independent of other components ($v_{1,t},v_{3,t},...,v_{n,t}$).
In practice, when there is a single noncausal state variable, the filtering algorithm of Section 4 can be completed by the two following steps:
step 7. The approximations of nonlinear causal innovations $\hat{v}_{2,t}(Z)$ can be computed from the estimated distributions as $\hat{v}_{2,t}(Z)=\Phi^{-1}(\hat{F}_{2,T}(\hat{Z}_{2,t}|\hat{Z}_{2,t-1}))$ by applying the formula of predictive density (5.6) with $l_{2}$ replaced by $\hat{l}_{2}$ and $g_{\eta_{2}}$ replaced by $\hat{g}_{\eta_{2}}$, i.e. the kernel-smoothed empirical density of $\hat{\eta}_{2,t},\,t=1,...,T$.
step 8. Next, $\hat{v}_{2,t}(Z)$ can be appended by the approximated nonlinear causal innovations $\hat{v}_{1,t}$, independent of $\hat{v}_{2,t}(Z)$, which are defined as: $\hat{v}_{1,t}=\Phi^{-1}(\hat{F}_{1|2,T}(\hat{Z}_{1,t}|\hat{Z}_{2,t},\hat{Z}_{t-1})),\;\mbox{for}\;n_{1}=1$.
@@ 6 Illustration
This Section illustrates the nonlinear forecasts from the mixed VAR(1) model and the nonlinear IRF analysis. Subsection 6.1 presents a simulation study that examines the oos forecasts from the model. In subsection 6.2, the semi-parametric GCov estimators, nonlinear innovations and IRFs are illustrated in an application to a bivariate series of US GDP rates and oil prices. Additional simulations and empirical results are provided in Online Appendices D and E.
@@ 6.1 Simulation Study
We consider a simulated bivariate mixed VAR(1) process with the following matrix of autoregressive coefficient: $\Phi=\left(\begin{array}[]{cc}0.7&-1.3\\ 0&2\end{array}\right),$ with eigenvalues 0.7 and 2, located inside and outside the unit circle, respectively . The errors follow a bivariate noise with independent components both t-student distributed with $\nu$ =4 degrees of freedom, mean zero, and variance equal to $\nu/(\nu-2)=2$. The matrix A is as follows: $A=\left(\begin{array}[]{cc}1&-1\\ 0&1\end{array}\right).$ The simulated paths of the series of length 500 is displayed in Figure 1. The solid (black) line represents process $(Y_{1t})$ and the dashed (red) line represents process $(Y_{2t})$. The sequence of spikes in the noncausal component $Y_{2t}=Z_{2t}$ impacts the component $Y_{1t}$ through the recursive form of matrix $\Phi$.
Let us now consider the forecasts based on parameter estimates. The Generalized Covariance (GCov) estimate of matrix $\Phi$ is obtained by minimizing the portmanteau statistic computed from the auto- and cross-correlations up to and including lag $H=10$ of the errors $\varepsilon_{t}=Y_{t}-\Phi_{1}Y_{t-1}$ and their squared values [see Gourieroux and Jasiak (2023)] [^10] The finite sample properties of the GCov estimator are illustrated in Gourieroux and Jasiak (2023).. The estimated autoregressive matrix is $\hat{\Phi}=\left(\begin{array}[]{cc}0.724&-1.452\\ -0.030&1.993\end{array}\right),$ with eigenvalues $\hat{\lambda}_{1}=0.690$, $\hat{\lambda}_{2}=2.027$, which are close to the true values $\lambda_{1}=J_{1}=0.7$ and $\lambda_{2}=J_{2}=2.0$. The standard errors of $\hat{\Phi}$ obtained by bootstrap are 0.023, 0.308, for the elements of the first row, and 0.009, 0.120 for the elements of the second row.
@@
After estimating $\Phi$, the GCov estimated errors $\hat{\varepsilon}_{t}=Y_{t}-\hat{\Phi}Y_{t-1}$ are computed. Matrices $A$ and $A^{-1}$ are identified from the real Jordan representation of matrix $\Phi$, up to scale factors. The estimated matrix $\hat{A}^{-1}$ computed from the normalized Jordan decomposition of $\hat{\Phi}$ is:
@@
$\hat{A}^{-1}=\left(\begin{array}[]{cc}0.022&0.025\\ -0.022&0.974\end{array}\right)$. It corresponds to the true matrix $A^{-1}=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)$ up to scale factors of about 0.02 and 0.97 for each column. We use these estimates to approximate the causal and noncausal components displayed in Figure 2.